# Transversal Hamilton cycles in digraph collections

### v1

@@ 4 Stable case
Thus, $|[n]\setminus\mathcal{C}_{\rm rem}|=|V\setminus U|\leq\lambda n+4\beta n<2\lambda n$ since $\beta\ll\lambda$, and $\delta^{0}(\mathcal{J})\geq(\frac{1}{2}-\mu-2\lambda)n\geq(\frac{1}{2}-3\lambda)n$.
Since $\lambda\ll\gamma,\alpha,\epsilon,\delta$ and $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, it is easy to see that $\mathcal{J}$ is $(\frac{\gamma}{2},\frac{\alpha}{2},2\epsilon,\frac{\delta}{2})$-stable. Apply Lemma 2.2 to $\mathcal{J}$ with parameters $(\epsilon_{0},1,d,L_{0})$. Let $\mathcal{R}$ be the reduced digraph collection of $\mathcal{J}$. Write $[L]$ for the common vertex set of digraphs in $\mathcal{R}$ where $L_{0}\leq L\leq n_{0}$, and $[M]$ for the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $U$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}_{\rm rem}$ and a digraph collection $\mathcal{J}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have that ${hi}\in R_{j}$ if and only if $\mathcal{J}_{hi,j}^{\prime}:=\{J_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq|U|\leq Mm+\epsilon_{0}n$ and $Mm\leq|\mathcal{C}_{\rm rem}|\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Based on Lemma 2.4, we obtain that, for each vertex $i\in[L]$, there are at least $(1-d^{1/4})M$ colors $j\in[M]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$; for each color $j\in[M]$, there are at least $(1-d^{1/4})L$ vertices $i\in[L]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$.
@@
Without loss of generality, assume that $P^{1},P^{2},\ldots,P^{s}$ are all the disjoint rainbow directed paths obtained in the above. Then $s\leq 2\sqrt{\epsilon_{0}}n+L<2^{-10}\beta n$. Let $x_{i},y_{i}$ be the start vertex and end vertex of $P^{i}$ for each $i\in[s]$ (where $x_{i}=y_{i}$ if $P^{i}$ has length 0). Clearly, $U=\bigcup_{i\in[s]}V(P^{i})$. It follows that $|\mathcal{C}_{\rm rem}\setminus{\rm col}(\bigcup_{i\in[s]}P^{i})|=s$.
Next, we do the following operation for each $i\in[s]$ in turn. Let $a_{i}\in\mathcal{C}_{\rm rem}$ be an arbitrary unused color. Choose an unused 4-tuple $Q_{i}:=(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})\in\mathcal{Q}$ where $j_{i}\in[r]$, $a_{i}\in L({v_{j_{i}}x_{i}})$ and $c_{j_{i}}^{\prime}\in L({y_{i}v_{j_{i}}^{\prime}})$. This is feasible since Claim 4.4 implies that there are at least $2^{-9}\beta n$ choices for $Q_{i}$, of which at most $s\leq 2^{-10}\beta n$ have been used. Now, $(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})$ is totally absorbable, so there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$ inside ${C}$. Then choose one of them, say $S_{i}:=x_{1}^{i}x_{2}^{i}x_{3}^{i}x_{4}^{i}$, whose vertices have not been previously chosen since $s<2^{-10}\beta n<2^{-10}\lambda^{2}n$, and whose colors are $b_{1}^{i},b_{2}^{i},b_{3}^{i}$ in order.
Notice that there remains $I\subseteq[r]$ such that the $(c_{j},c_{j}^{\prime},v_{j},v_{j}^{\prime})\in\mathcal{Q}$ with $j\in I$ are precisely the $4$-tuples which were not chosen to be some $Q_{i}$. For each one, there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j}$-absorbing paths of $(v_{j},v_{j})$ and disjoint rainbow directed $c_{j}^{\prime}$-absorbing paths of $(v_{j}^{\prime},v_{j}^{\prime})$ inside $C$. Since $\beta<\lambda^{2}$, there are two such rainbow directed rainbow paths $T_{i},T_{i}^{\prime}$ for each $(c_{j},v_{j})$ and $(c_{j}^{\prime},v_{j}^{\prime})$, which are disjoint and whose vertices have not previously been chosen.
@@
For each $i\in[s]$, we replace $S_{i}$ by $x_{1}^{i}x_{2}^{i}v_{j_{i}}x_{i}P^{i}y_{i}v_{j_{i}}^{\prime}x_{3}^{i}x_{4}^{i}$ with colors $b_{1}^{i},c_{j_{i}},a_{i}$ followed by the colors inherited from $P^{i}$, followed by $c_{j_{i}}^{\prime},b_{2}^{i},b_{3}^{i}$. That is, we have replaced $S_{i}$ by a directed rainbow path with the same endpoints, vertices $V(S_{i})\cup V(P^{i})\cup\{v_{j_{i}},v_{j_{i}}^{\prime}\}$, and colors ${\rm col}(S_{i})\cup{\rm col}(P^{i})\cup\{c_{j_{i}},c_{j_{i}}^{\prime},a_{i}\}$. For each $i\in I$, we replace $T_{i}:=y_{1}^{i}y_{2}^{i}y_{3}^{i}y_{4}^{i}$ by $y_{1}^{i}y_{2}^{i}v_{i}y_{3}^{i}y_{4}^{i}$ where colors are inherited except ${\rm col}(y_{2}^{i}v_{i})=c_{i}$ and ${\rm col}(v_{i}y_{3}^{i})={\rm col}(y_{2}^{i}y_{3}^{i})$. We do a similar replacement of $T_{i}^{\prime}$, using new vertex $v_{i}^{\prime}$ and new color $c_{i}^{\prime}$. So we have replaced $T_{i}$ (resp. $T_{i}^{\prime}$) by a rainbow directed path with the same endpoints, vertices $V(T_{i})\cup\{v_{i}\}$ (resp. $V(T_{i}^{\prime})\cup\{v_{i}^{\prime}\}$) and colors ${\rm col}(T_{i})\cup\{c_{i}\}$ (resp. ${\rm col}(T_{i}^{\prime})\cup\{c_{i}^{\prime}\}$). Thus we have obtained a directed rainbow cycle using vertices $V(C)\cup\bigcup_{i\in[s]}V(P^{i})\cup\{v_{i},v_{i}^{\prime}:i\in[r]\}=V$ and colors ${\rm col}(C)\cup\bigcup_{i\in[s]}{\rm col}(P^{i})\cup\{c_{i},c_{i}^{\prime}:i\in[r]\}=[n]$ where each color is used at most once (and hence exactly once). That is, $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired. ∎
@@ 5 Extremal case
Choose constants $\mu,\alpha,\gamma,\epsilon,\delta$ such that
$0<\frac{1}{n}\ll\mu\ll\alpha\ll\gamma,\epsilon\ll\delta\ll 1\ \text{and}\ \delta\ll\min\left\{\zeta,\frac{1}{2}-\zeta-\epsilon\right\},$
where $\zeta$ is defined in Lemma 3.3. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\left(\frac{1}{2}-\mu\right)n$. Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. By Lemma 4.3, we know $\mathcal{D}$ is not $(\gamma,\alpha)$-strongly stable. Hence there are at least $(1-\gamma)n$ digraphs in $\mathcal{D}$ such that each of them is $\alpha$-extremal. Without loss of generality, assume that $D_{i}$ is $\alpha$-extremal for each $i\in[(1-\gamma)n]$. Note that every $\alpha$-extremal digraph is also $\epsilon$-extremal, since $\alpha\ll\epsilon$. Denote
@@
[…]
Hence $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|+|\tilde{\mathcal{C}}_{3}|\geq(1-\gamma)n$. By Lemma 3.3, we know that $D_{i}$ has a characteristic partition $(\tilde{A_{i}},\tilde{B_{i}},\tilde{L_{i}})$ for each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, or a characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each $i\in\tilde{\mathcal{C}}_{3}$.
Under the above characteristic partition, applying Lemma 4.3 yields that $\mathcal{D}$ is not $(\epsilon,\delta)$-weakly stable. For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{1}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For $i\in\tilde{\mathcal{C}}_{3}$, define $I^{2}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{3}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then we have
@@
  1. (B1)
    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta\tilde{A}_{i}|<\delta n$ and $|\tilde{B}_{1}\Delta\tilde{B}_{i}|<\delta n$ for all $i\in(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$,
  2. (B2)
    if $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels, for all $k\in[4]$ and all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{2}$ we have $|W_{1}^{k}\Delta W_{i}^{k}|<\delta n$, in particular $|C_{1}^{k}\Delta C_{i}^{k}|<2\delta n$,
  3. (B3)
    if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\Delta W_{i}^{1}|<\delta n$ and either $|\tilde{A}_{1}\Delta W_{i}^{2}|<\delta n$ or $|\tilde{A}_{1}\Delta W_{i}^{4}|<\delta n$ for all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$.
Note that if ${\bf(B3)}$ holds, then for each $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$, there is an $i_{0}\in\{2,4\}$ such that
$2\delta n>|\tilde{A}_{1}\Delta W_{i}^{1}|+|\tilde{A}_{1}\Delta W_{i}^{i_{0}}|\geq|W_{i}^{1}\Delta W_{i}^{i_{0}}|>2\delta n,$
a contradiction. This implies that exactly one of $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$ holds. Based on (B1)-(B2), we proceed our proof by considering the following two theorems.
@@
Assume that $\delta$ is defined in (5). Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $|\tilde{\mathcal{C}}_{3}|<2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.

### v2

@@ 4 Stable case
Thus, $\big|[n]\setminus\mathcal{C}_{\rm rem}\big|=|V\setminus U|\leq\lambda n+4\beta n<2\lambda n$ since $\beta\ll\lambda$, and $\delta^{0}(\mathcal{J})\geq(\frac{1}{2}-\mu-2\lambda)n\geq(\frac{1}{2}-3\lambda)n$.
Since $\lambda\ll\gamma,\alpha,\epsilon,\delta$ and $\mathcal{D}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, it is easy to see that $\mathcal{J}$ is $(\frac{\gamma}{2},\frac{\alpha}{2},2\epsilon,\frac{\delta}{2})$-stable. Applying Lemma 2.2 to $\mathcal{J}$ with parameters $(\epsilon_{0},1,d,L_{0})$, we obtain a reduced digraph collection $\mathcal{R}$. Denote by $[L]$ the common vertex set of the digraphs in $\mathcal{R}$, where $L_{0}\leq L\leq n_{0}$, and by $[M]$ the set of color clusters. Thus there is a partition $V_{0},V_{1},\ldots,V_{L}$ of $U$ and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}_{\rm rem}$ and a digraph collection $\mathcal{J}^{\prime}$ satisfying (i)-(v) of Lemma 2.2. Therefore, for $\{(h,i),j\}\in{[L]\choose 2}\times[M]$, we have that ${hi}\in R_{j}$ if and only if $\mathcal{J}_{hi,j}^{\prime}:=\{J_{c}^{\prime}[V_{h},V_{i}]:c\in\mathcal{C}_{j}\}$ is $(\epsilon_{0},d)$-regular. Furthermore, $Lm\leq|U|\leq Mm+\epsilon_{0}n$ and $Mm\leq|\mathcal{C}_{\rm rem}|\leq Lm+\epsilon_{0}n$, so $|L-M|\leq\frac{\epsilon_{0}n}{m}\leq\frac{\epsilon_{0}L}{1-\epsilon_{0}}$. Thus, we may assume that $M=L$ at the expense of assuming the slightly worse bound $|V_{0}|+|\mathcal{C}_{0}|\leq 3\epsilon_{0}n$. Based on Lemma 2.4, we obtain that, for each vertex $i\in[L]$, there are at least $(1-d^{1/4})M$ colors $j\in[M]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$; for each color $j\in[M]$, there are at least $(1-d^{1/4})L$ vertices $i\in[L]$ such that $d_{R_{j}}^{+}(i),d_{R_{j}}^{-}(i)\geq(\frac{1}{2}-4\lambda)L$.
@@
Without loss of generality, assume that $P^{1},P^{2},\ldots,P^{s}$ are all the disjoint rainbow directed paths obtained in the above. Then $s\leq 2\sqrt{\epsilon_{0}}n+L<2^{-10}\beta n$. Let $x_{i},y_{i}$ be the start vertex and end vertex of $P^{i}$ for each $i\in[s]$ (where $x_{i}=y_{i}$ if $P^{i}$ has length 0). Clearly, $U=\bigcup_{i\in[s]}V(P^{i})$. It follows that $\big|\mathcal{C}_{\rm rem}\setminus{\rm col}(\bigcup_{i\in[s]}P^{i})\big|=s$.
Next, we do the following operation for each $i\in[s]$ in turn. Let $a_{i}\in\mathcal{C}_{\rm rem}$ be an arbitrary unused color. Choose an unused 4-tuple $Q_{i}:=(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})\in\mathcal{Q}$ where $j_{i}\in[r]$, $a_{i}\in L({v_{j_{i}}x_{i}})$ and $c_{j_{i}}^{\prime}\in L({y_{i}v_{j_{i}}^{\prime}})$. This is feasible since Claim 4.4 implies that there are at least $2^{-9}\beta n$ choices for $Q_{i}$, of which at most $s\leq 2^{-10}\beta n$ have been used. Now, $(c_{j_{i}},c_{j_{i}}^{\prime},v_{j_{i}},v_{j_{i}}^{\prime})$ is totally absorbable, so there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j_{i}}$-absorbing path of $(v_{j_{i}},v_{j_{i}}^{\prime})$ inside C{C}. Then choose one of them, say $S_{i}:=x_{1}^{i}x_{2}^{i}x_{3}^{i}x_{4}^{i}$, whose vertices have not been previously chosen since $s<2^{-10}\beta n<2^{-10}\lambda^{2}n$, and whose colors are $b_{1}^{i},b_{2}^{i},b_{3}^{i}$ in order.
Notice that there remains $I\subseteq[r]$ such that the $(c_{j},c_{j}^{\prime},v_{j},v_{j}^{\prime})\in\mathcal{Q}$ with $j\in I$ are precisely the $4$-tuples which were not chosen to be some $Q_{i}$. For each one, there are at least $\lambda^{2}n$ disjoint rainbow directed $c_{j}$-absorbing paths of $(v_{j},v_{j})$ and disjoint rainbow directed $c_{j}^{\prime}$-absorbing paths of $(v_{j}^{\prime},v_{j}^{\prime})$ inside $C$. Since $\beta<\lambda^{2}$, there are two such rainbow directed paths $T_{i},T_{i}^{\prime}$ for each $(c_{j},v_{j})$ and $(c_{j}^{\prime},v_{j}^{\prime})$, which are disjoint and whose vertices have not previously been chosen.
@@
For each $i\in[s]$, we replace $S_{i}$ by $x_{1}^{i}x_{2}^{i}v_{j_{i}}x_{i}P^{i}y_{i}v_{j_{i}}^{\prime}x_{3}^{i}x_{4}^{i}$ with colors $b_{1}^{i},c_{j_{i}},a_{i}$ followed by the colors inherited from $P^{i}$, followed by $c_{j_{i}}^{\prime},b_{2}^{i},b_{3}^{i}$. That is, we have replaced $S_{i}$ by a rainbow directed path with the same endpoints, vertices $V(S_{i})\cup V(P^{i})\cup\{v_{j_{i}},v_{j_{i}}^{\prime}\}$, and colors ${\rm col}(S_{i})\cup{\rm col}(P^{i})\cup\{c_{j_{i}},c_{j_{i}}^{\prime},a_{i}\}$. For each $i\in I$, we replace $T_{i}:=y_{1}^{i}y_{2}^{i}y_{3}^{i}y_{4}^{i}$ by $y_{1}^{i}y_{2}^{i}v_{i}y_{3}^{i}y_{4}^{i}$ where colors are inherited except ${\rm col}(y_{2}^{i}v_{i})=c_{i}$ and ${\rm col}(v_{i}y_{3}^{i})={\rm col}(y_{2}^{i}y_{3}^{i})$. We do a similar replacement of $T_{i}^{\prime}$, using new vertex $v_{i}^{\prime}$ and new color $c_{i}^{\prime}$. So we have replaced $T_{i}$ (resp. $T_{i}^{\prime}$) by a rainbow directed path with the same endpoints, vertices $V(T_{i})\cup\{v_{i}\}$ (resp. $V(T_{i}^{\prime})\cup\{v_{i}^{\prime}\}$) and colors ${\rm col}(T_{i})\cup\{c_{i}\}$ (resp. ${\rm col}(T_{i}^{\prime})\cup\{c_{i}^{\prime}\}$). Thus we have obtained a rainbow directed cycle using vertices $V(C)\cup\bigcup_{i\in[s]}V(P^{i})\cup\{v_{i},v_{i}^{\prime}:i\in[r]\}=V$ and colors ${\rm col}(C)\cup\bigcup_{i\in[s]}{\rm col}(P^{i})\cup\{c_{i},c_{i}^{\prime}:i\in[r]\}=[n]$ where each color is used at most once (and hence exactly once). That is, $\mathcal{D}$ contains a transversal directed Hamilton cycle, as desired. ∎
@@ 5 Extremal case
In this section, we consider the extremal case, in which most digraphs in the collection are $(\epsilon,\mathrm{ECk})$-extremal for some $k\in[3]$. The following result follows directly from Lemma 4.3.
Lemma 5.1.
Assume that $0<\frac{1}{n}\ll 8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$ and $m=(1-4\sqrt{\delta})n$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $\mathcal{D}$ contains no transversal directed Hamilton cycles, then by swapping indices if necessary, one of the following holds:
  1. (B1)
    either $D_{i}$ is $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal and it admits an $\epsilon$-characteristic partition $(\tilde{A}_{i},\tilde{B}_{i},\tilde{L}_{i})$ for each $i\in[m]$,
  2. (B2)
    or $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal and it admits an $\epsilon$-characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ for each $i\in[m]$.
In (B1), for every $i\in[m]$, $D_{i}$ also admits a partition $(A_{i},B_{i},L_{i})$ satisfying
  • (C1)
    if $i=1$ then $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$; for $i\geq 2$,
  • (C2)
    if $D_{i}$ is $(\epsilon,K_{\lceil\frac{n}{2}\rceil}\cup K_{\lfloor\frac{n}{2}\rfloor})$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{D_{i}}(v,Y_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
  • (C3)
    if $D_{i}$ is $(\epsilon,K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor})$-extremal, then for $\{Y,Z\}=\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{D_{i}}(v,Z_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
  • (C4)
    subject to (C2)-(C3), the partition $(A_{i},B_{i},L_{i})$ is chosen such that $|A_{i}\cup B_{i}|$ is maximized for each $i\in[2,m]$.
Furthermore, for every $i\in[m]$ and $Y\in\{A,B\}$, we have
$|L_{i}|\leq 2\epsilon n\ \ \text{and}\ \ |{Y}_{1}{\triangle}Y_{i}|<2\delta n.$
Proof.
Choose constants $\alpha,\gamma,\epsilon,\delta$ such that
$0<\frac{1}{n}\ll\alpha\ll\gamma,8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1.$
Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a digraph collection on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. By Lemma 4.3, we know that $\mathcal{D}$ is not $(\gamma,\alpha)$-strongly stable. Hence, there are at least $(1-\gamma)n$ digraphs in $\mathcal{D}$ such that each of them is $\alpha$-extremal. Without loss of generality, assume that $D_{i}$ is $\alpha$-extremal for each $i\in[(1-\gamma)n]$. Note that every $\alpha$-extremal digraph is also $\epsilon$-extremal, since $\alpha<\epsilon$. Denote
$\tilde{\mathcal{C}}_{1}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}1)\text{-extremal}\},$
$\tilde{\mathcal{C}}_{2}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}2)\text{-extremal}\},$
$\tilde{\mathcal{C}}_{3}$ $:=\{i\in[n]:D_{i}\ \text{is}\ (\epsilon,{\rm EC}3)\text{-extremal}\}.$
@@
Under the above characteristic partition, applying Lemma 4.3 again yields that $\mathcal{D}$ is not $(\epsilon,\delta)$-weakly stable. For each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{1}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For each $i\in\tilde{\mathcal{C}}_{3}$, define $I^{2}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. For each $i\in\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2}$, define $I^{3}_{i}$ to be the set of $j\in\tilde{\mathcal{C}}_{3}$ such that $D_{i}$ and $D_{j}$ are $\delta$-crossing. Then we have
@@
  • if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\triangle\tilde{A}_{i}|<\delta n$ and $|\tilde{B}_{1}\triangle\tilde{B}_{i}|<\delta n$ for all $i\in(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$,
  • if $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels, for all $k\in[4]$ and all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{2}$ we have $|W_{1}^{k}\triangle W_{i}^{k}|<\delta n$, in particular $|C_{1}^{k}\triangle C_{i}^{k}|<2\delta n$,
  • if $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then by swapping labels one has $|\tilde{A}_{1}\triangle W_{i}^{1}|<\delta n$ and either $|\tilde{A}_{1}\triangle W_{i}^{2}|<\delta n$ or $|\tilde{A}_{1}\triangle W_{i}^{4}|<\delta n$ for all $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$.
Note that if last itemize holds, then for each $i\in\tilde{\mathcal{C}}_{3}\setminus I^{3}$, there is an $i_{0}\in\{2,4\}$ such that
$2\delta n>|\tilde{A}_{1}\triangle W_{i}^{1}|+|\tilde{A}_{1}\triangle W_{i}^{i_{0}}|\geq|W_{i}^{1}\triangle W_{i}^{i_{0}}|>2\delta n,$
a contradiction. This implies that exactly one of $|\tilde{\mathcal{C}}_{1}|+|\tilde{\mathcal{C}}_{2}|\geq 2\sqrt{\delta}n$ and $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$ holds.
[…]
  • if $D_{i}$ is $(\epsilon,{\rm EC}1)$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{i}^{+}(v,Y_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
  • if $D_{i}$ is $(\epsilon,{\rm EC}2)$-extremal, then for $Y\in\{A,B\}$ we have $\tilde{Y_{i}}\subseteq Y_{i}$ and $d_{i}^{+}(v,Z_{i})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for each vertex $v\in Y_{i}$,
  • subject to the above two conditions, the partition $(A_{i},B_{i},L_{i})$ is chosen such that $|A_{i}\cup B_{i}|$ is maximized for each $i\in[2,m]$.
Denote $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$. Hence for every $i\in[m]$ and $Y\in\{A,B\}$, we have $|Y_{1}\triangle Y_{i}|\leq|\tilde{Y_{1}}\triangle\tilde{Y_{i}}|+|\tilde{L}_{i}|\leq\delta n+2{\epsilon}n<2\delta n$, as desired. ∎
Based on Lemma 5.1, we proceed with the proof by considering the following two theorems.
@@
Assume that $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1$ and $m=(1-4\sqrt{\delta})n$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $D_{i}$ is $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal for each $i\in[m]$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.